# ANID: How Far Are We? Evaluating the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance

### v1

@@ Abstract
In the rapidly evolving field of Artificial Intelligence Generated Content (AIGC), one of the key challenges is distinguishing AI-synthesized images from natural images. Despite the remarkable capabilities of advanced AI generative models in producing visually compelling images, significant discrepancies remain when these images are compared to natural ones. To systematically investigate and quantify these discrepancies, we introduce an AI-Natural Image Discrepancy Evaluation benchmark aimed at addressing the critical question: how far are AI-generated images (AIGIs) from truly realistic images? We have constructed a large-scale multimodal dataset, the Distinguishing Natural and AI-generated Images (DNAI) dataset, which includes over 440,000 AIGI samples generated by 8 representative models using both unimodal and multimodal prompts, such as Text-to-Image (T2I), Image-to-Image (I2I), and Text vs. Image-to-Image (TI2I). Our fine-grained assessment framework provides a comprehensive evaluation of the DNAI dataset across five key dimensions: naive visual feature quality, semantic alignment in multimodal generation, aesthetic appeal, downstream task applicability, and coordinated human validation. Extensive evaluation results highlight significant discrepancies across these dimensions, underscoring the necessity of aligning quantitative metrics with human judgment to achieve a holistic understanding of AI-generated image quality. Code is available at https://github.com/ryliu68/ANID.
Introduction
@@
With the rapid advancement of deep learning techniques, the proliferation of Artificial Intelligence Generated Content (AIGC) has garnered significant attention across various domains, such as e-commerce, gaming, medicine, animation, and autonomous driving (Li et al. 2024; Qian et al. 2024). AI-generated Images (AIGI) have been one of the mainstream forms of AIGC, and a variety of AI image generative models have been proposed to make the generated synthetic images as realistic as natural images, ranging from the earlier generative adversarial networks (GANs) (Tao et al. 2023, 2022), advanced diffusion models (DMs) (Xu et al. 2023b; Wei et al. 2023) to the large multimodal generative model like DALL·E (Ramesh et al. 2022).
@@
Despite the proliferation of AI image generative models, AI-generated images are not qualified for real-world applications, and the discrepancies between generated synthetic images and realistic natural images still exist (Li et al. 2023a). Therefore, many AI-generated image quality assessment and evaluation methods have been proposed (Wang et al. 2023; Xu et al. 2023a; Hu et al. 2023). For example, Wang et al. (Wang et al. 2023) established an AIGC Image Quality Assessment Database called AIGCIQA2023 from 6 text-to-image generative models, and devised a subjective evaluation framework to assess human visual preferences for each AI-generated image from three aspects, e.g., quality, authenticity, and text-to-image correspondence. PKU-I2IQA (Yuan et al. 2023) constructed a human perception-based image-to-image database named PKU-I2IQA and conducted a subject analysis based on both no-reference and full-reference methods. Different from prior studies focusing on establishing the AIGI dataset and conducting a subjective evaluation with human feedback, another group of studies (e.g. Pick-a-Pic (Kirstain et al. 2023), HPS v2 (Wu et al. 2023), TIFA (Hu et al. 2023), and ImageReward (Xu et al. 2023a)) focused on training a unified score model as the automatic evaluation metrics for AIGC image quality assessment (AIGIQA) measures, in contrast with traditional image quality assessment for natural images measures like Inception Score (Salimans et al. 2016) and FID (Heusel et al. 2017).
However, some issues and problems have not yet been fully answered and solved. First, prior studies only considered unimodal promopted generation, e.g. Text-to-Image (T2I) or Image-to-Image (I2I), ignoring multimodal prompted content like Text vs.Image-to-Image (TI2I). The data size of the established T2I or I2I dataset is relatively small (around thousands of images), and might be insufficient for comprehensive analysis and evaluation. Second, prior studies typically focused on the perceptual quality of AIGC images with traditional image quality assessment metrics, the text-to-image correspondence, and subjective human preference. No studies have fully answered the important questions about what discrepancies exist between generated synthetic images and realistic natural images, and they lacked a systematic and compressive assessment and evaluation to investigate and interpret the discrepancies between AI-generated synthetic images and realistic natural images. The investigation and understanding of the discrepancies are key ways to realize the practical application of AIGC images in real-world scenarios and make breakthroughs in AIGC.
@@
To address the above challenges, we proposed an AI-Natural Image Discrepancy Evaluation Benchmark to investigate and interpret what discrepancies still remained between AI-generated images and natural images, finally answering the important question “How far are AI generative models with respect to the visual forms of images?”. Especially, we have two important contributions with reference to the key problems. Large Multimodal Evaluation Dataset construction: we have curated a large multimodal evaluation dataset called DNAI (Distinguishing Natural and AI-generated) dataset, with about 440,000 AIGI generated from 8 representative generative models based on both unimodal and mutimodal guidance including Text-to-Image (T2I), Image-to-Image (I2I), and Text vs. Image-to-Image (TI2I). The data size of our DNAI dataset scales to 100x that of prior datasets. Fine-grained Evaluation Framework, we propose a fine-grained evaluation framework to conduct the systematic and comprehensive assessment and evaluation of DNAI, covering 5 diverse and important aspects, including naive visual feature quality, semantic alignment among multimodal generation, aesthetic appeal, downstream applicability, and the coordinated human validation.
@@
Through our DNAI dataset and fine-grained evaluation framework, we conduct extensive benchmark analysis and evaluation and conclude key insights to answer the discrepancy questions:
@@
  • Significant discrepancies in key Areas: AI-generated images exhibit substantial discrepancies from natural images in terms of quantitative measures from all aspects, with about 10% to 30%.
@@
  • Multimodal Alignment: Different prompted generation might have different semantic alignment scores. Generated images prompted with texts including both T2I and TI2I demonstrated remarkable semantic alignments.
@@
  • Downstream Task Applicability: Significant differences exist in the usability of AI-generated images vs. natural ones in downstream tasks, highlighting the need for further improvements of generative models to ensure practical applicability in real-world scenarios.
@@
  • Human Evaluation vs. Quantitative Metrics: Human evaluation results reveal larger discrepancies compared with quantitative metrics. It validates the necessity of incorporating human evaluation for coordination.
Related Work
@@
Over the past decades, numerous image quality assessment methods have been developed to evaluate traditional natural images (N. et al. 2015). These methods have focused on various visual features and properties, such as perceptual appearance (Zhang et al. 2018), naturalness (Ma et al. 2018), and aesthetics (Esfandarani and Milanfar 2018). For instance, BRISQUE evaluates natural image quality by analyzing spatial domain features, while PIQE assesses perceptual quality through block-based image segmentation, both serving as effective no-reference metrics. Other notable measures include FID and Inception Scores, which assess the quality and diversity of generated images, and SSIM (Wang et al. 2004) and PSNR, which quantify structural similarity and pixel-level accuracy. Beyond quantitative measures, human evaluations have also been leveraged to estimate natural image quality. For example, NIMA (Esfandarani and Milanfar 2018), proposed by Talebi and Milanfar, uses a deep CNN trained on human-rated images to predict aesthetic quality. Similarly, Wong et al. introduced the AVA dataset, which facilitates aesthetic assessment based on human preferences (Murray, Marchesotti, and Perronnin 2012).
@@
Recently, AI-generated content (AIGC) has seen significant advancements with the rise of generative models. Several methods and benchmarks have been introduced to evaluate AI-generated images (Zhang et al. 2023; Hu et al. 2023). These methods primarily focus on three aspects: perceptual quality, text-image correspondence, and aesthetics (Xu et al. 2023a; Wang et al. 2023), often training unified models to assess the overall quality of AI-generated images. For example, ImageReward (Xu et al. 2023a) presents a framework for aligning image generation with human preferences, AGIQA-3k (Li et al. 2023a) offers a comprehensive benchmark for assessing AI-generated image quality, and QBench (Wu et al. 2024) evaluates content quality across multiple dimensions.
@@
Despite advancements, little research has delved into the fine-grained differences between AI-generated and natural images, and existing datasets are often too small for comprehensive evaluation. To address these gaps, we propose an AI-Natural Image Difference Evaluation Benchmark, featuring a large Distinguishing AI-Natural Image Dataset and a systematic, fine-grained evaluation framework to thoroughly explore these differences.
Figure 1: Overview of AI-Natural Image Discrepancy Evaluation Benchmark (ANID): (a) A large Distinguishing AI-Natural Image (DANI) dataset is curated, guided by three types of prompts. (b) A fine-grained evaluation framework is proposed to measure and evaluate the potential discrepancies between AI-generated images and natural images from five aspects covering naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and coordinated human assessment.
AI-Natural Image Discrepancy Evaluation Benchmark
To tackle the challenges of current AI-generated image evaluation, we construct an AI-Natural Image Discrepancy evaluation benchmark (ANID) to evaluate the potential discrepancies between AI-generated images and natural images. Figure 1 shows the overview of our ANID benchmark.
@@
Our ANID benchmark contributes to two core components: (1) A Distinguishing AI-Natural Image (DANI) Dataset is constructed, in which we collect AI-generated images from diverse generative models for real natural images in MS COCO dataset (Lin et al. 2014) based on three types of guidance prompts e.g. text-only, image-only and text-image prompts; (2) We devise a systematic and comprehensive evaluation framework to measure and evaluate the potential differences between AI-generated images and real natural images from five aspects including naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and the coordinated human assessment. In the following parts, we introduce the DANI Dataset and the evaluation framework in detail.
Distinguishing AI-Natural Image Dataset
@@
For each text-image sample in the referenced natural image set, we collect AI-generated images from 8 representative generative models, guided by three types of prompts, including text-only prompts, image-only prompts, and text-image prompts. The 8 representative generative models are from diverse generative models ranging from the earlier generative adversarial networks (GANs) (e.g. GALIP (Tao et al. 2023), DF-GAN (Tao et al. 2022)) to the recent benchmarking diffusion models (DMs) like Stable Diffusion v1.4, v1.5, v2.1, XL (Rombach et al. 2022), Versatile-Diffusion (Xu et al. 2023b) and OpenAI DALL·E 2 (Ramesh et al. 2022). The referenced natural images and the generated images comprise our distinguishing AI-Natural image dataset. More details about data construction and collection are presented in Section Appendix A.
Table 1 shows the statistics and properties of our DANI Dataset compared with the existing evaluation dataset. Our DANI Dataset has a total of 440,000 AI-generated images from 8 representative generated models. For details, please refer to Appendix Sec. A. Compared with state-of-the-art evaluation datasets, our DANI Dataset has convincing contributions:
  • DANI Dataset is the most extensive dataset for AI-generated image evaluation, and the data size scales to 100x.
  • DANI Dataset is guided by both unimodal prompts and multi-modal prompts (text-image). Prior datasets barely considered text prompts.
  • Diverse generative models are exploited, ranging from the earlier GAN models, the recent popular diffusion models, and the influential large commercial generative model DALL·E 2 issued by OpenAI.
Our large and comprehensive dataset works as the foundation of the evaluation benchmark and enables thorough evaluation and analysis of the differences between AI-generated images and realistic natural images.
@@
DNAI (Ours) 8 25,000 5,000 25,000 440,000
Table 1: Comparison of DANI dataset and Counterparts.
Fine-Grained Evaluation Framework
@@
On top of the DANI dataset, we construct a fine-grained evaluation framework to systematically measure and interpret the differences between AI-generated images and realistic natural images from five different aspects, including naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and coordinated human assessment.
1. Naive Image Quality:
@@
We leverage traditional image quality assessment methods for natural images to measure the AI-Natural image differences. In order to perform fine-grained analysis and exploration, we estimate the AI-Natural image differences from three levels by considering the inherent properties of image-style content, ranging from the low-level visual features, frame-level visual features to the holistic content distribution.
@@
  • First, we use the pixel-level similarity measures like SSIM, LPIPS, and DISTS to compute the visual similarities between the paired AI-generated image and natural image and use the inverse similarity values to quantify the AI-Natural image differences from the low-level visual view.
@@
  • Second, we use frame-level visual features to qualify the perceptional quality of images and compare the global-level visual quality differences between AI-generated images and natural images from the high-level visual view. The exploited frame-level visual features include PIQE (N. et al. 2015), IL-NIQE (Zhang, Zhang, and Bovik 2015), MUSIQ (Ke et al. 2021), DBCNN (Zhang et al. 2020), LIQE (Ma et al. 2018), Inception Score (Salimans et al. 2016), CLIPIQA (Radford et al. 2021a), TReS (Gu et al. 2015), HyperIQA (Su et al. 2020), UNIQUE (Zhang et al. 2021), BRISQUE , NIQE (Mittal, Soundararajan, and Bovik 2013), NRQM (Ma et al. 2018), etc. These metrics provide insights into global-level visual quality differences.
@@
  • Third, we investigate the visual content distribution to capture the holistic differences between AI-generated images and natural images by computing the FID and Inception Score of the whole AI-generated image dataset and the referenced natural image set.
@@
Based on the different levels of image quality measures, our framework can evaluate the fine-grained AI-Natural image differences, covering both low-level and high-level. More details about the utilized traditional image quality assessment measures are described in Appendix Sec. C.
2. Semantic Alignment:

### v2

# D-Judge: How Far Are We? Accessing the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance

@@ Abstract
In Artificial Intelligence Generated Content (AIGC), distinguishing AI-synthesized images from natural ones remains a key challenge. Despite advancements in generative models, significant discrepancies persist. To systematically investigate and quantify these discrepancies, we introduce an AI-Natural Image Discrepancy accessing benchmark (D-Judge) aimed at addressing the critical question: how far are AI-generated images (AIGIs) from truly realistic images? We construct D-ANI, a dataset with 5,000 natural images and over 440,000 AIGIs generated by nine models using Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I) prompts. Our framework evaluates the discrepancy across five dimensions: naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and human validation. Results reveal notable gaps, emphasizing the importance of aligning metrics with human judgment. Source code and datasets are available at https://shorturl.at/l83W2.
Artificial Intelligence Generated Content, AI-synthesized Images, Multimodal, Discrepancy, Evaluation Framework
1 Introduction
@@
With the rapid advancement of deep learning techniques, the proliferation of Artificial Intelligence Generated Content (AIGC) has garnered significant attention across various domains, such as e-commerce, gaming, medicine, animation, and autonomous driving (Li et al., 2024; Qian et al., 2024). AI-generated Images (AIGI) have been one of the mainstream forms of AIGC, and a variety of AI image generative models have been proposed to make the generated synthetic images as realistic as natural images, ranging from the earlier generative adversarial networks (GANs) (Tao et al., 2023, 2022), advanced diffusion models (DMs) (Xu et al., 2023b; Wei et al., 2023) to the large multimodal generative model like DALL·E (Ramesh et al., 2022).
Figure 1: Toy example demonstrating the key aspects of D-Judge.
@@
Despite the rapid development of AI image generative models, AI-generated images still fall short of real-world application standards due to discrepancies with realistic natural images (Li et al., 2023a). To address this, various AI-generated image quality assessment methods have emerged (Wang et al., 2023; Xu et al., 2023a; Hu et al., 2023). For instance, Wang et al. (Wang et al., 2023) introduced the AIGCIQA2023 database and a subjective evaluation framework based on quality, authenticity, and text-image correspondence. Similarly, PKU-I2IQA (Yuan et al., 2023) developed a perception-based database for image-to-image generation using no-reference and full-reference methods. Other studies, such as Pick-a-Pic (Kirstain et al., 2023), HPS v2 (Wu et al., 2023), TIFA (Hu et al., 2023), and ImageReward (Xu et al., 2023a), focused on training unified score models for automatic AIGC image quality assessment, moving beyond traditional metrics like Inception Score (Salimans et al., 2016) and FID (Heusel et al., 2017).
Despite advancements in AIGC research, several critical issues remain unresolved. First, there is a lack of systematic and comprehensive studies addressing the discrepancies between AI-generated images and realistic natural images, which is crucial for enabling the practical application of AIGC images in real-world scenarios and achieving meaningful breakthroughs in the field. Most prior studies primarily focus on perceptual quality using traditional image quality metrics, text-to-image correspondence, or subjective human preferences, leaving a gap in understanding the broader implications of these discrepancies. Second, previous research predominantly considers unimodal prompts, such as T2I or I2I, while neglecting multimodal prompts like TI2I, and the datasets used are often small, typically comprising only a few thousand images, which limits their capacity for comprehensive analysis and evaluation.
@@
To address the above challenges, we proposed an AI-Natural Image Discrepancy Evaluation Benchmark (D-Judge) to investigate and interpret what discrepancies still remained between AI-generated images and natural images, finally answering the important question “How far are AI generative models with respect to the visual forms of images?”. Especially, we have two important contributions with reference to the key problems. Large Multimodal Evaluation Dataset construction: we have curated a large multimodal evaluation dataset called D-ANI (Distinguishing Natural and AI-generated) dataset, with about 440,000 AIGI generated from 9 representative generative models based on both unimodal and mutimodal guidance including Text-to-Image (T2I), Image-to-Image (I2I), and Text and Image-to-Image (TI2I). The data size of our DNAI dataset scales to 100x that of prior datasets. Fine-grained Evaluation Framework, we propose a fine-grained evaluation framework to conduct the systematic and comprehensive assessment and evaluation of DNAI, covering 5 diverse and important aspects, including naive visual feature quality, semantic alignment among multimodal generation, aesthetic appeal, downstream applicability, and the coordinated human validation. The toy example is illustrated in Figure 1. For example, given a text prompt, an image is generated using a generative model and paired with a natural reference image retrieved from the natural database. D-Judge then evaluates the discrepancy rate between the two images across the aforementioned dimensions.
@@
Through our D-ANI dataset and fine-grained evaluation framework D-Judge, we conduct extensive benchmark analysis and evaluation and conclude key insights to answer the discrepancy questions:
@@
  • Significant discrepancies in key Areas: AIGIs exhibit substantial discrepancies from natural images in terms of quantitative measures from all aspects, up to 40%, 33.57%, 27.66%, and 96.95% in frame-level image quality, semantic alignment, aesthetic appeal, and downstream applicability, respectively.
@@
  • Multimodal Alignment: Different prompted generations might have different semantic alignment discrepancies compared to natural ones. Generated images prompted with texts including both T2I and TI2I demonstrated less semantic discrepancy than I2I.
@@
  • Downstream Task Applicability: Significant differences in usability between AI-generated and natural images are observed in downstream tasks, particularly in fine-grained object recognition and VQA, with discrepancy rates reaching as high as 94.29% and 96.95%, respectively. This underscores the need for further advancements in generative models to enhance their practical applicability in real-world scenarios.
@@
  • Human Evaluation vs. Quantitative Metrics: Human evaluation results reveal larger discrepancies compared with quantitative ones. It validates the necessity of incorporating human evaluation for coordination.
2 Related Work
@@
Over the past decades, numerous image quality assessment methods have been developed to evaluate natural images (N. et al., 2015). These methods have focused on various visual features and properties, such as perceptual appearance (Zhang et al., 2018), naturalness (Ma et al., 2018), and aesthetics (Esfandarani & Milanfar, 2018). For instance, BRISQUE evaluates natural image quality by analyzing spatial domain features, while PIQE assesses perceptual quality through block-based image segmentation, both serving as effective no-reference metrics. Other notable measures include FID and Inception Scores, which assess the quality and diversity of generated images, and SSIM (Wang et al., 2004) and PSNR, which quantify structural similarity and pixel-level accuracy. Beyond quantitative measures, human evaluations have also been leveraged to estimate natural image quality. For example, NIMA (Esfandarani & Milanfar, 2018), proposed by Talebi and Milanfar, uses a deep CNN trained on human-rated images to predict aesthetic quality. Similarly, Wong et al. introduced the AVA dataset, which facilitates aesthetic assessment based on human preferences (Murray et al., 2012).
@@
Recently, AI-generated content (AIGC) has seen significant advancements with the rise of generative models. Several methods and benchmarks have been introduced to evaluate AI-generated images (Zhang et al., 2023; Hu et al., 2023). These methods primarily focus on three aspects: perceptual quality, text-image correspondence, and aesthetics (Xu et al., 2023a; Wang et al., 2023), often training unified models to assess the overall quality of AI-generated images. For example, ImageReward (Xu et al., 2023a) presents a framework for aligning image generation with human preferences, AGIQA-3k (Li et al., 2023a) offers a comprehensive benchmark for assessing AI-generated image quality, and QBench (Wu et al., 2024a) evaluates content quality across multiple dimensions.
@@
Despite advancements, little research has delved into the fine-grained differences between AI-generated and natural images, and existing datasets are often too small for comprehensive evaluation. To address these gaps, we propose D-Judge as AI-Natural image difference Assessment Benchmark, featuring a large Distinguishing AI-Natural Image Dataset and a systematic, fine-grained evaluation framework to thoroughly explore these differences.
3 D-Judge
Figure 2: The D-ANI dataset of D-Judge, guided by three types of prompts. Each generated image is paired with both the corresponding text and image from the naturally collected dataset (e.g., COCO).
Figure 3: The framework of D-Judge evaluates the Discrepancy Rate ($DR$) between AI-generated images and natural reference images across five dimensions: ➊ Naive Quality, ➋ Alignment, ➌ Aesthetic, and ➍ Applicability, complemented by ➎ human validation. The images with green borders represent reference natural images, while those with yellow borders are AI-generated counterparts. This comprehensive evaluation integrates pixel-level, frame-level, and distribution-level quality checks, alongside advanced tools like OpenAI CLIP, OpenAI Laion Aesthetic Evaluator, Qalign, ViT, U2Net, and BLIP2 to robustly assess the gap between the two image types. Where $CS$ means CLIP Score for the text and given images, and $AES$ means Aesthetic score.
@@ 3 D-Judge
To tackle the challenges of current AI-generated image evaluation, we construct an AI-Natural Image Discrepancy evaluation benchmark (D-Judge) to assess the potential discrepancies between AI-generated images and natural images and answer how far AIGIs are from Natural Images. Firstly, a Distinguishing AI-Natural Image (D-ANI) Dataset is constructed, in which we collect AI-generated images from diverse generative models for real natural images in MS COCO dataset (Lin et al., 2014) based on three types of guidance prompts, e.g., text-only, image-only and text-image prompts, as shown in Figure 2; Then, we devise a systematic and comprehensive evaluation framework as Figure 3 illustrated to measure and evaluate the potential differences between AI-generated images and real natural images from five aspects including naive image quality, semantic alignment, aesthetic appeal, downstream applicability, and the coordinated human assessment. In the following parts, we introduce the D-ANI Dataset and the evaluation framework in detail.
3.1 Distinguishing AI-Natural Image Dataset
@@
For each text-image sample in the referenced natural image set, we collect AI-generated images from 9 representative generative models, guided by three types of prompts, including text-only prompts, image-only prompts, and text-image prompts. The 9 representative generative models are from diverse generative models ranging from the earlier generative adversarial networks (GANs) (e.g. GALIP (Tao et al., 2023), DF-GAN (Tao et al., 2022)) to the recent diffusion models (DMs) like Stable Diffusion v1.4 (SD14), v1.5 (SD15), v2.1 (SD21), XL (SDxl) (Rombach et al., 2022), Versatile-Diffusion (VD) (Xu et al., 2023b), OpenAI DALL·E 2 (D·E 2) (Ramesh et al., 2022) and DALL·E 3 (D·E 3) (OpenAI, 2024). The referenced natural images and the generated images comprise our distinguishing AI-Natural image dataset. More details about data construction and collection are presented in Appendix Sec. B.
Table 1 shows the statistics and properties of our D-ANI Dataset compared with the existing evaluation dataset. Our D-ANI Dataset has a total of 440,000 AI-generated images from 9 representative generated models. For details, please refer to Appendix Sec. B. Compared with state-of-the-art evaluation datasets, our D-ANI Dataset has convincing contributions:
  • D-ANI Dataset is the most extensive dataset for AI-generated image evaluation, and the data size scales to 100x.
  • D-ANI Dataset is guided by both unimodal prompts and multi-modal prompts (text-image). Prior datasets barely considered text prompts.
  • Comprehensive and Diverse generative models are exploited, ranging from the earlier GAN models, the recent popular diffusion models, and the influential large commercial generative model DALL·E 2 and DALL·E 3 issued by OpenAI.
Our large and comprehensive dataset works as the foundation of the assessment benchmark and enables thorough evaluation and analysis of the differences between AI-generated images and realistic natural images.
Table 1: Comparison of D-ANI dataset and Counterparts.
@@
D-ANI (Ours) 9 25,000 5,000 25,000 445,000
3.2 D-Judge Framework
@@
On top of the D-ANI dataset, we construct the fine-grained D-Judge to systematically measure and interpret the differences between AI-generated images and realistic natural images from five key different aspects as following aspects.
1. Naive Image Quality Discrepancy:
@@
We leverage traditional image quality assessment methods for natural images to measure the AI-Natural image differences. In order to perform fine-grained analysis and exploration, we estimate the AI-Natural image differences from three levels by considering the inherent properties of image-style content, ranging from the pixel-level visual features, frame-level visual features to the holistic content distribution.
@@
  • First, we use the full-referenced pixel-level similarity measures like SSIM, LPIPS, and DISTS to compute the visual similarities between the paired AI-generated image and natural image and use these calculated values to quantify the AI-Natural image differences from the low-level visual view.
@@
  • Second, we use non-referenced frame-level visual features to qualify the perceptional quality of images and compute the frame-level visual quality difference rates between AI-generated images and natural images from the high-level visual view. The exploited frame-level visual features include PIQE (N. et al., 2015), IL-NIQE (Zhang et al., 2015), MUSIQ (Ke et al., 2021), DBCNN (Zhang et al., 2020), LIQE (Ma et al., 2018), Inception Score (Salimans et al., 2016), CLIPIQA (Radford et al., 2021b), TReS (Gu et al., 2015), HyperIQA (Su et al., 2020), UNIQUE (Zhang et al., 2021), BRISQUE , NIQE (Mittal et al., 2013), NRQM (Ma et al., 2018), etc. The Discrepancy Rate ($DR$) of these metrics provides insights into frame-level visual quality differences.
@@
  • Third, we investigate the full-reference visual content distribution to capture the holistic differences between AI-generated images and natural images by computing the FID, CLIP-FID, and Inception Score of the AI-generated image set and the referenced natural image set. The results of these set-level measures can reflect the difference between AIGIs and natural ones from an integral perspective.
@@
Based on the different levels of image quality measures, our framework can assess and measure the fine-grained AI-Natural images differences with Full-reference and Non-reference angles, covering both low-level and high-level. More details about the utilized traditional image quality assessment measures are described in Appendix Sec. C.
2. Semantic Alignment Discrepancy: